OpenStudy (anonymous):

A country's population, in millions, is given by the exponential model P = 78.2e0.047t, where t represents years after 2000. After how many years will the country have a population of 210 million? (Round answer to the nearest tenth.) A) 12.0 years B) 21.0 years C) 29.8 years D) 37.8 years

OpenStudy (anonymous):

so, we need to solve for Time t \[P(t) = 78.2* e^{0.047 t}\] Now, let P(t) = 210 millions: \[210 = 78.2* e^{0.047 t}\] ; divide both sides by 78.2 you get: \[2.685242 = e^{0.047 t}\]; now take the natural log ( ln ) of both sides: \[ln 2.685242 = ln e^{0.047 t}\] ; which is equal to: \[ln 2.685242 = {0.047 t}\] because ln e = 1, and so : t = 21.017 years. therefore the answer is B) 21 years. anything you don't understand let me know!

OpenStudy (anonymous):

wow thank you i used to be good at this but i completely forgot it.